OpenStudy (anonymous):

find domain and ?range of two variables function f(x,y)=square root of y-x ?

OpenStudy (anonymous):

plzzzzzzzzz ans me

OpenStudy (anonymous):

think about the restrictions on x and y for their difference to be a positive number

OpenStudy (anonymous):

^he's right

OpenStudy (anonymous):

sory i din't under stand

OpenStudy (anonymous):

domain is the set of values of x for which function is defined... which is \[\left\{ \left( x,y \right)\epsilon R ^{2} : y \ge x \right\}\]

OpenStudy (anonymous):

how would you make y-x=positive number

OpenStudy (anonymous):

y=x

OpenStudy (anonymous):

lets see how....if y<x in your function, you'll get a negative number in the square root...which is undesirable

OpenStudy (anonymous):

if we take squre on both side m r rite ?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

so?

OpenStudy (anonymous):

do you know what is DOMAIN?

OpenStudy (anonymous):

these are two variable function

OpenStudy (anonymous):

`

OpenStudy (anonymous):

yea...A domain is the set of numbers which are valid values for the independent variables

OpenStudy (anonymous):

mens imaginar are not in domain ?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

when is y-x negative?

OpenStudy (anonymous):

indepent nd dependent what is difference b/w both

OpenStudy (anonymous):

<0 then negative ? m r rite

OpenStudy (anonymous):

???????????

OpenStudy (anonymous):

yea so domain is 'those values of y and x' for which y-x> or equal to 0 (non-negative)... solve y-x>or equal to 0 y>or equal to x

OpenStudy (anonymous):

are we clear on domain?

OpenStudy (anonymous):

now clear buddy

OpenStudy (anonymous):

Good

OpenStudy (anonymous):

about range ?

OpenStudy (anonymous):

now the range is the set of value of function you get when you put values you got in domain...

OpenStudy (anonymous):

If you put y> or equal to x in you function do we get positive f(x,y) or negative?

OpenStudy (anonymous):

positive

OpenStudy (anonymous):

coz < 0r = sign

OpenStudy (anonymous):

right...positive..so you range will be \[\left\{ f \in R : f \ge 0 \right\}\]

OpenStudy (anonymous):

ok thanxxxxxxxxxxxxxxxxxxxxx alotttttttttttt

OpenStudy (anonymous):

aha welcomio

OpenStudy (anonymous):

:)